OpenStudy (anonymous):

Multiply: 2 the square root of 11 • 8 the square root of 22 8 the square root of 11 10 the square root of 33 64 the square root of 22 176 the square root of 2

OpenStudy (anonymous):

\[2 \times \sqrt 11 \times 8 \sqrt 22 =2 \times 8 \times \sqrt {11 \times 22} =16 \sqrt {11 \times 11 \times 2}\] \[ =16 \times 11\sqrt { 2} =176 \sqrt { 2}\] Hence ' 176 the square root of 2' is the correct option.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Welcome